area of rectangle is 28 sq ft. one side is 2ft longer. what are dimensions?
Answer by checkley77(12844)

L=W+2
2L+2W=28
2(W+2)+2W=28
2W+4+2W=28
4W=28-4
4W=24
W=24/4
W=6 FT. IS THE SHORTER SIDE.
L=6+2=8 FT. IS THE LONGER SIDE.
PROOF:
2*6+2*8=28
12+16=28
28=28